# Lossless Model Compression via Joint Low-Rank Factorization Optimization

### v1

@@ 1 Introduction
Figure 2: Weight error and loss values of different models, where Rank1$>$Rank2$>$Rank3$>$Rank4. When the weight error $\delta$ is the lowest, the model loss $L$ does not reach the minimum value.
@@ 2 Background
[…]
@@
where $f(\cdot)$ represents the loss of the model on a dataset, $\mathbf{E}$ stands for expectation, $m$ is the size of the dataset, $\ell(\cdot)$ is the loss function for a sample, and $(x_{i},y_{i})$ denotes a sample in the dataset along with its corresponding label, $L(\cdot)$ represents the loss function, such as the cross-entropy function; $h_{i}$, with $i\in[1,...,n]$ represents the ($n-i+1$)th layer in the neural network, $W=(w_{n}^{T},w_{n-1}^{T},\cdots,w_{1}^{T})^{T}$, where $w_{i}$ is the parameter in $h_{i}(\cdot)$, and for the reason of a unified format, $h_{n+1}$ denotes the sample $x$.
@@
where $b$ is the bias, $Y$ is the output of the linear layer, and the factorization of $W$ is obtained from Eq. 1. With Eq. 2, we can store and compute $L$ and $R$ instead of $W$. The total number of parameters of $L$ and $R$ is $Nk+Mk$. The reduced parameters and computation are $NM-(Nk+Mk)$. When the weight matrix $M=N$, the rank $k$ is less than $0.5M$ and the model size will be reduced. Similarly, for singular value decomposition, $W$ is approximated as $\widehat{W}=USV^{T}$, where $U$ and $V$ are orthonormal, and $S$ is diagonal.
@@ 3 Lossless Joint Factorization Strategy
¯$\bar{\ell}_{k}(w,x_{j},y_{j})=L(h_{1}(h_{2}(\cdots h_{n}(x_{i},w_{n}+\delta^{k}_{n})\cdots,$
@@
¯$\bar{\ell}_{k}(w,x_{i},y_{i})-\ell(w,x_{i},y_{i})=\sum_{i=1}^{n}\frac{\partial\ell_{k}}{\partial w_{i}}\cdot\delta^{k}_{i},$ (6)
@@
$\operatorname*{min}_{\delta^{k}\in\Delta}\bar{f}(w)-f(w)$ $=\frac{1}{m}\sum_{i=1}^{n}\sum_{(x_{j},y_{j})\in\mathbb{D}}\frac{\partial\ell}{\partial w_{i}}\cdot\delta^{k}_{i},$ (9)
@@
1:Neural network $M$, lossless rank list $A$.
@@
9:         Initialize $L_{c}$, $R_{c}$triangleright Temporary matrix
@@
13:                 {$L_{c}$, $R_{c}$, $Loss$} append to list $A$
@@
18:    Search for the minimized Loss in the List $A$.
19:    $W_{a}$ = $L_{a}$ $R_{a}$triangleright Final decomposition matrix
@@ 4 Lossless and Compact Optimization
9:         Initialize $L_{c}$, $R_{c}$triangleright Temporary matrix
@@
13:                 $W_{a}$ = $L_{a}$ $R_{a}$triangleright Final matrix
@@
where ^$\hat{\mathcal{R}}(\mathcal{A})$ represents the emperical error, and $\lambda_{min}$ represents the minimum eign-value of the loss Hermitian matrix. For the loss Hermitian matrix [13], since the model has been trained, $\lambda_{min}\approx 0$. Based on the above equation, we can observe that when the effective rank ratio $p$ is low and the sparsity (1 – $p$) is relatively high, the generalization error decreases with increasing sparsity. However, as the effective rank decreases and the sparsity increases, there comes a point where the number of trainable parameters is much lower than the number required to represent the distribution of the dataset, resulting in underfitting. Therefore, there is an optimal effective rank that allows the weights to be decomposed. This ensures that the algorithm is effective.
@@ 5.1 Datasets and Details
+ ACIQ [1] 77.140 93.382 0.940222 -- 73.00%
+ Ours(L) 77.008 93.420 0.920947 -- 68.00%
+ Ours(C) 76.284 93.114 0.938988 -- 78.00%
@@ 5.3 Ablation
         Layer          $\epsilon$          $\Delta$ $Loss$
@@
Theory to Practice Mapping and Constraints. We focus on whether the noise introduced by practical low-rank factorization satisfies our differential neighborhoods theory analysis. Our differential theory conditions that the neighborhoods need to be sufficiently small, and this corresponds to the performance implications of the actual factorization. We factorize the transformer layer. From Table 7, the factorization noise of the model’s layers with different ranks leads to different upper bounds on the noise, and the variation in losses $\Delta$ $Loss$ due to this noise is very small. This satisfies our analysis of differential neighborhoods. When too much noise is introduced, the change in loss $\Delta$ $Loss$ is too large, which is also beyond the analytical scope of mathematical differentiation. Low rank introduces noise well beyond our lossless constraints. As shown in Figure 3, when the rank is too low, the noise introduced by factorization is too large, and the lossless restriction conditions in the optimization objective fail, causing the loss to rise beyond the original model. This makes our algorithm incapable of lossless factorization at current very low ranks.

### v2

# Loss-Minimizing Model Compression via Joint Factorization Optimization

@@ 1 Introduction
Figure 2: Weight error and loss values of different models, where Rank1>>Rank2>>Rank3>>Rank4. When the weight error $\delta$ is the lowest, the model loss $L$ does not reach the minimum value.
@@ 2 Background
$\begin{aligned} \min_{W}f(W)=\mathbf{E}_{Sample}\ell(W,Sample)=\frac{1}{m}\sum_{(x_{i},y_{i})\in\mathbb{D}}\ell(W,x_{i},y_{i})\end{aligned},$ (1)
@@
where $f(\cdot)$ represents the loss of the model on a dataset, $\mathbf{E}$ stands for expectation, $m$ is the size of the dataset, $\ell(\cdot)$ is the loss function for a sample, and $(x_{i},y_{i})$ denotes a sample in the dataset along with its corresponding label, $L(\cdot)$ represents the loss function, such as the cross-entropy function; $h_{i}$, with $i\in[1,...,n]$ represents the (ni+1n-i+1)th layer in the neural network, $W=(w_{n}^{T},w_{n-1}^{T},\cdots,w_{1}^{T})^{T}$, where $w_{i}$ is the parameter in $h_{i}(\cdot)$, and for the reason of a unified format, $h_{n+1}$ denotes the sample $x$.
@@
where $b$ is the bias, $Y$ is the output of the linear layer, and the factorization of $W$ is obtained from Eq. 1. With Eq. 2, we can store and compute $L$ and $R$ instead of $W$. The total number of parameters of $L$ and $R$ is $Nk+Mk$. The reduced parameters and computation are NM(Nk+Mk)NM-(Nk+Mk). When the weight matrix $M=N$, the rank $k$ is less than $0.5M$ and the model size will be reduced. Similarly, for singular value decomposition, $W$ is approximated as $\widehat{W}=USV^{T}$, where $U$ and $V$ are orthonormal, and $S$ is diagonal.
@@ 3 Lossless Joint Factorization Strategy
$\bar{\ell}_{k}(w,x_{j},y_{j})=L(h_{1}(h_{2}(\cdots h_{n}(x_{i},w_{n}+\delta^{k}_{n})\cdots,$
@@
$\bar{\ell}_{k}(w,x_{i},y_{i})-\ell(w,x_{i},y_{i})=\sum_{i=1}^{n}\frac{\partial\ell_{k}}{\partial w_{i}}\cdot\delta^{k}_{i},$ (6)
@@
$\operatorname*{min}_{\delta^{k}\in\Delta}\bar{f}(w)-f(w)$ $=\frac{1}{m}\sum_{i=1}^{n}\sum_{(x_{j},y_{j})\in\mathbb{D}}\frac{\partial\ell}{\partial w_{i}}\cdot\delta^{k}_{i},\ \$(9)
@@
1:Neural network $M$, lossless rank list AA.
@@
9:    Initialize $L_{c}$, $R_{c}$ $\triangleright$ Temporary matrix
@@
13:      $L_{c}$, $R_{c}$, LossLoss} append to list AA
@@
18:  Search for the minimized Loss in the List AA.
19:  $W_{a}$ = $L_{a}$ $R_{a}$ $\triangleright$ Final decomposition matrix
@@ 4 Lossless and Compact Optimization
9:    Initialize $L_{c}$, $R_{c}$ $\triangleright$ Temporary matrix
@@
13:      $W_{a}$ = $L_{a}$ $R_{a}$ $\triangleright$ Final matrix
@@
where $\hat{\mathcal{R}}(\mathcal{A})$ represents the emperical error, and $\lambda_{min}$ represents the minimum eign-value of the loss Hermitian matrix. For the loss Hermitian matrix [13], since the model has been trained, $\lambda_{min}\approx 0$. Based on the above equation, we can observe that when the effective rank ratio $p$ is low and the sparsity (1 – $p$) is relatively high, the generalization error decreases with increasing sparsity. However, as the effective rank decreases and the sparsity increases, there comes a point where the number of trainable parameters is much lower than the number required to represent the distribution of the dataset, resulting in underfitting. Therefore, there is an optimal effective rank that allows the weights to be decomposed. This ensures that the algorithm is effective.
@@ 5.1 Datasets and Details
+ ACIQ [1] 77.140 93.382 0.940222 - 73.00%
+ Ours(L) 77.008 93.420 0.920947 - 68.00%
+ Ours(C) 76.284 93.114 0.938988 - 78.00%
@@ 5.3 Ablation
   Layer    $\epsilon$    $\Delta$ LossLoss
@@
Theory to Practice Mapping and Constraints. We focus on whether the noise introduced by practical low-rank factorization satisfies our differential neighborhoods theory analysis. Our differential theory conditions that the neighborhoods need to be sufficiently small, and this corresponds to the performance implications of the actual factorization. We factorize the transformer layer. From Table 7, the factorization noise of the model’s layers with different ranks leads to different upper bounds on the noise, and the variation in losses $\Delta$ LossLoss due to this noise is very small. This satisfies our analysis of differential neighborhoods. When too much noise is introduced, the change in loss $\Delta$ LossLoss is too large, which is also beyond the analytical scope of mathematical differentiation. Low rank introduces noise well beyond our lossless constraints. As shown in Figure 3, when the rank is too low, the noise introduced by factorization is too large, and the lossless restriction conditions in the optimization objective fail, causing the loss to rise beyond the original model. This makes our algorithm incapable of lossless factorization at current very low ranks.